### v2

@@ Appendix A Appendix
1$\Rightarrow$2) Create a decreasing chain of conditioning events $(C^{1},...,C^{n})$ as follows. Let $C^{1}=S_{-i}$. For each $k>1$, let $C^{k}$ be the union of all the $s_{-i}\in S_{-i}$ such that $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$. Let $n$ be the smallest $k$ such that $C^{k+1}=\emptyset$. Consider the list of measures $\hat{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C^{1}),...,\bar{\mu}_{i}(\cdot|C^{n}))$. For each $h_{i}\in H_{i}$, let $k$ be the smallest $j\leq n$ such that $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})>0$ — it exists by $C^{n+1}=\emptyset$. By $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})=0$ for every $j<k$, it follows that $S_{-i}(h_{i})\subseteq C^{j}$. Hence, $\mu_{i}(\cdot|h_{i})$ can be derived from $\bar{\mu}_{i}(\cdot|C^{k})$ by updating, and $\bar{\mu}_{i}(\cdot|C^{k})$ is also the first measure in $\hat{\mu}_{i}$ that gives positive probability to $S_{-i}(h_{i})$. There remains to show that $\hat{\mu}_{i}$ is a LCPS. For each $s_{-i}\in S_{-i}$, by $C^{n+1}=\emptyset$, $\bar{\mu}_{i}(s_{-i}|C^{k})>0$ for some $k\leq n$, so $\hat{\mu}_{i}$ has full joint support. Moreover, by $s_{-i}\in C^{k}$, $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$, and by $\bar{\mu}_{i}(s_{-i}|C^{k})>0$, $s_{-i}\not\in C^{j}$ for every $j>k$, which implies $\bar{\mu}_{i}(s_{-i}|C^{j})=0$. Hence, $\hat{\mu}_{i}$ has disjoint supports and is therefore a LCPS. $\blacksquare$
Proposition 2 implies that the main result of this paper also applies to belief consistency as it was defined for sequential games. With this, the characterization of complete consistency with the coherence rule among discounted odds ratios can be rewritten using traditional odds ratios in place of discounted odds ratios, because in this context the two coincide by Proposition 1.
@@
Siniscalchi (2020) (the working paper of Siniscalchi 2022) comes to a similar characterization of complete consistency starting from a different idea. Consider a belief system that is isomorphic to a conditional probability system defined over the collection of observable events. Such belief system thus satisfies the chain rule between contingencies that correspond to nested observations, but the chain rule may have little bite when the observable events are not nested, and the corresponding beliefs are not disciplined by the preliminary observation of a larger event. Therefore, Siniscalchi (2020) generalizes the chain rule to apply also between overlapping, but non-nested observable events. We conclude this section by showing the convergence between Siniscalchi’s approach and mine.
Call “generalized self-odds ratio” a generalized odds ratio between $s_{-i}$ and itself.
@@
All the generalized odds ratios of each pair are identical if and only if all generalized self-odds ratios are $1$.
Proof. Only if. Fix a generalized self-odds ratios. Break it into two and invert one of the two halves. We have two concatenations of odds ratios that start and end with the same element. Thus, the two corresponding generalized odds ratios are identical by assumption. Therefore, the generalized self-odds ratio is $1$.
If. Fix two concatenations of odds ratios where either no odds ratio is infinite or there is no zero. If both contain a zero, the corresponding generalized odds ratios are both zero. If they both contain an infinite, they are both infinite. If one contains an infinite and the other contains a zero, the generalized self-odds ratio obtained by inverting one of the two is zero or infinite, which violates the assumption. Otherwise, the generalized self-odds ratio (in one direction) is finite. By assumption, it is $1$. Therefore the two generalized odds ratios are identical. $\blacksquare$
@@
A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if for every $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$,
@@
If part. Fix $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$. If both sides of equation 11 are zero, consistency holds. Otherwise, suppose without loss of generality that the right-hand side is not zero, so that equation 11 can be rewritten as
@@
Note that equation 12 holds if for every $s_{-i}\in E$ such that the ratio $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ is not indeterminate,
@@
because this means that all the non-indeterminate ratios $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ are identical, which implies that the ratio $\mu^{i}(E|h_{i}^{1})/\mu^{i}(E|h_{i}^{n})$ takes the same value too. We are going to show that for every $m=2,...,n$, there is $s_{-i}^{m}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that
@@
then, we can rewrite equation 13 as
@@
which is true by complete consistency (cf. Remark 1). Let $s_{-i}^{m}$ be any $s_{-i}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that $\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})>0$; one exists by the fact that the right-hand side of 11 is not zero. By complete consistency, for each $s_{-i}^{\prime}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$, we have
@@
Only if part. By Remark 1 we can show that for each $s_{-i},\in S_{-i}$, for every generalized self-odds ratio, equation 14 holds. Without loss of generality, we can assume that the generalized self-odds ratio is not infinite (because an infinite one can be inverted and becomes zero), so that no denominator in equation 14 is zero. For each $m=2,...,n$, by consistency, equation 11 for $E=\left\{s_{-i}^{m}\right\}$ and $\left(h_{i}^{1},...,h_{i}^{n}\right)=(h_{i}^{m-1},h_{i}^{m})$ yields
@@
Then, substituing in 14, we obtain equation 13, which coincides with equation 12 for $E=\left\{s_{-i}\right\}$, and thus holds by consistency.
@@ A.2 Completion of the proofs of the theorems
Proof of Theorem 1: $1\Rightarrow 2$
Fix a non-terminal contingency $h$ and let $h^{1},...,h^{n}$ be the next contingencies. Let $\alpha^{1},...,\alpha^{n}$ denote the weights that verify the definition of Bayes-plausibility. We construct probability distributions $(p(h^{m}|h,s))_{m=1}^{n}$ of reaching each $h^{m}$ given $h$ and each $s\in S(h)$. Fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, let $p(h^{m}|h,s)=0$ for all $s\in\mathcal{S}$. Suppose now that $\alpha^{m}>0$. For each $s\in\mathcal{S}$, let
$p(h^{m}|h,s)$ $=$ $\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\text{ \ \ if }s\in\mathrm{supp}\mu(\cdot|h),$
$p(h^{m}|h,s)$ $=$ $\alpha^{m}\text{ \ \ \ \ \ \ \ \ \ if }s\not\in\mathrm{supp}\mu(\cdot|h).$
@@
Now fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, then
$\sum_{s\in S(h)}p(h^{m}|h,s)=0.$
If $\alpha^{m}>0$, by Bayes-plausibility
@@
$\frac{p(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}p(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=\frac{\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\mu(s|h)}{\alpha^{m}\sum_{s^{\prime}\in\mathrm{supp}\mu(\cdot|h)}\mu(s|h^{m})}=\mu(s|h^{m}),\qquad\forall s\in\mathcal{S}$
where the last equality follows from (15). Thus, $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(p(h^{m}|h,s))_{s\in S(h)}$ whenever possible.
Proof of Theorem 2: $3\Rightarrow 1$
We prove the counterpositive. Suppose that there exist two different GORs between some $s$ and $s^{\prime}$. Then, inverting all the DORs in the second GOR and multiplying by the first GOR, [^15] Since the two GORs are different, they cannot be both infinite or both zero, therefore after inverting one of the two their product is not indeterminate. we obtain a new GOR between $s$ and itself, which is not $1$; write it as
$r:=o(s^{0},s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{n}|h^{n}),$
where $n\geq 2$, $s^{0}=s^{n}=s$, and for each $k=1,...,n$,
$o(s^{k-1},s^{k}|h^{k})=\frac{p(h^{k}|s^{k-1})}{p(h^{k}|s^{k})}\frac{\mu(s^{k-1}|h^{k})}{\mu(s^{k}|h^{k})}.$ (16)
Wwithout loss of generality, we can suppose that:
$s^{0},...,s^{n-1}$ are all distinct, because if $s^{j}=s^{k}$, either $r^{\prime}=o(s^{j},s^{j+1}|h^{j})\cdot...\cdot o(s^{k-1},s|h^{k})$, or $r^{\prime\prime}=r/r^{\prime}$ is a new GOR between $s$ and itself that is not $1$.
$r<1$, because if a GOR is larger than $1$, inverting all the DORs we obtain a GOR smaller than $1$ (still between $s$ and itself).
With this, we construct a Dutch Book that the agent is willing to accept. For every $h\in\mathcal{H}\backslash\left\{h^{1},...,h^{n}\right\}$, let $g(\widetilde{s}|h)=0$ for every $\widetilde{s}\in\mathcal{S}$, so that $i$ is willing to accept $g(\cdot|h)$. Fix $\varepsilon>0$. Let $g(\cdot|h^{1})$ be the following gamble on $S(h^{1})$:
@@
$g(s^{1}|h^{1})$ $=$ $-g(s^{0}|h^{1})\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon=\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon,$ (17)
and $g(s^{\prime}|h^{1})=0$ for every $s^{\prime}\not=s^{0},s^{1}$. Recursively, for each $m=2,...,n$, let $g(\cdot|h^{m})$ be the following gamble on $S(h^{m})$:
$g(s^{m-1}|h^{m})$ $=$ $-g(s^{m-1}|h^{m-1})\frac{p(h^{m-1}|s^{m-1})}{p(h^{m}|s^{m-1})}-\varepsilon,$ (18)
$g(s^{m}|h^{m})$ $=$ $-g(s^{m-1}|h^{m})\frac{\mu(s^{m-1}|h^{m})}{\mu(s^{m}|h^{m})}+\varepsilon,$ (19)
@@
and $g(s^{\prime}|h^{m})=0$ for every $s^{\prime}\not=s^{m},s^{m-1}$. For each $m=1,...,n$, the agent is willing to accept $g(\cdot|h^{m})$ because
$\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})+\mu(s^{m}|h^{m})\cdot g(s^{m}|h^{m})$
$=$ $\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})-g(s^{m-1}|h^{m})\cdot\mu(s^{m-1}|h^{m})+\varepsilon\cdot\mu(s^{m}|h^{m})$
$=$ $\varepsilon\cdot\mu(s^{m}|h^{m})>0\text{,}$
where $\mu(s^{m}|h^{m})>0$ because $r\not=\infty$. Thus, $i$ is willing to accept $g=(g(\cdot|h))_{h\in\mathcal{H}}$. If $m\not=n$, since $s^{m}\not=s^{m^{\prime}}$ for every $m^{\prime}\not=m$, we also have
$\sum_{h\in\mathcal{H}}p(h|s^{m})g(s^{m}|h)$ $=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})+p(h^{m+1}|s^{m})g(s^{m}|h^{m+1})$
$=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})-p(h^{m+1}|s^{m})g(s^{m}|h^{m})\frac{p(h^{m}|s^{m})}{p(h^{m+1}|s^{m})}-p(h^{m+1}|s^{m})\varepsilon$
$=$ $-p(h^{m+1}|s^{m})\varepsilon<0,$
where $p(h^{m+1}|s^{m})>0$ because $s^{m}\in S(h^{m+1})$ (DORs at a contingency are only defined between consistent states). Thus, $s^{m}$ verifies the strict inequality condition of a Dutch book. Every $\widetilde{s}\in\mathcal{S}\backslash\left\{s^{1},...,s^{n}\right\}$ trivially satisfies the weak inequality condition. Finally, we need to show that, for some $\varepsilon>0$, $s$ satisfies the weak inequality condition as well, so that $g$ is a Dutch book. Note that, for $\varepsilon=0$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)$
$=$ $p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{n}|s)g(s^{n-1}|h^{n})\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
$=$ $-p(h^{1}|s)+p(h^{n}|s)g(s^{n-1}|h^{n-1})\frac{p(h^{n-1}|s^{n-1})}{p(h^{n}|s^{n-1})}\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-1}|h^{n-1})o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-1})\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-2})\frac{p(h^{n-2}|s^{n-2})}{p(h^{n-1}|s^{n-2})}\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)+p(h^{n-2}|s^{n-2})g(s^{n-2}|h^{n-2})o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{1}|s^{1})g(s^{0}|h^{1})\frac{\mu(s|h^{1})}{\mu(s^{1}|h^{1})}\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h)$
$=$ $-p(h^{1}|s)+p(h^{1}|s)r<0\text{,}$

### v3

@@ Appendix A Appendix
Proposition 3 implies that the Dutch-book theorem also applies to belief consistency as it was defined for sequential games. With this, the characterization of complete consistency with the coherence rule among discounted odds ratios can be rewritten using traditional odds ratios in place of discounted odds ratios, because in this context the two coincide by Proposition 2.
@@
Siniscalchi (2020) (the working paper of Siniscalchi 2022) comes to a similar characterization of complete consistency starting from a different idea. Consider a belief system that is isomorphic to a conditional probability system defined over the collection of observable events. Such belief system thus satisfies the chain rule between contingencies that correspond to nested observations, but the chain rule may have little bite when the observable events are not nested, and the corresponding beliefs are not disciplined by the preliminary observation of a larger event. Therefore, Siniscalchi (2020) generalizes the chain rule to apply also between overlapping, but non-nested observable events. We conclude this section by showing the convergence between Siniscalchi’s approach and ours.
Call “simple generalized self-odds ratio” a simple generalized odds ratio between $s_{-i}$ and itself.
@@
All the simple generalized odds ratios of each pair are identical if and only if all simple generalized self-odds ratios are $1$.
Proof.
Only if Fix a simple generalized self-odds ratio. Break it into two and invert one of the two halves. We have two concatenations of odds ratios that start and end with the same element. Thus, the two corresponding simple generalized odds ratios are identical by assumption. Therefore, the simple generalized self-odds ratio is $1$.
If Fix a pair and two concatenations of odds ratios of this pair where either no odds ratio is infinite or there is no zero. If both contain a zero, the corresponding simple generalized odds ratios are both zero. If they both contain an infinite, they are both infinite. If one contains an infinite and the other contains a zero, the simple generalized self-odds ratio obtained by inverting one of the two is zero or infinite, which violates the assumption. Otherwise, the simple generalized self-odds ratio (in one direction) is finite. By assumption, it is $1$. Therefore the two simple generalized odds ratios are identical. $\blacksquare$
@@
A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if for every $(h_{i}^{1},...,h_{i}^{n})\in H_{i}^{n}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$,
@@
If Fix $(h_{i}^{1},...,h_{i}^{n})\in H_{i}^{n}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$. If both sides of equation (12) are zero, consistency holds. Otherwise, suppose without loss of generality that the right-hand side is not zero, so that equation (12) can be rewritten as
@@
Note that equation (13) holds if for every $s_{-i}\in E$ such that the ratio $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ is not indeterminate,
@@
because this means that all the non-indeterminate ratios $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ are identical, which implies that the ratio $\mu^{i}(E|h_{i}^{1})/\mu^{i}(E|h_{i}^{n})$ takes the same value too. We are going to show that for every $m\in\left\{2,...,n\right\}$, there is $s_{-i}^{m}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that
@@
then, we can rewrite equation (14) as
@@
which is true by complete consistency (cf. Remark 1). Let $s_{-i}^{m}$ be any $s_{-i}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that $\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})>0$; one exists by the fact that the right-hand side of (12) is not zero. By complete consistency and Corollary 1, for each $s_{-i}^{\prime}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$, we have
@@
Only if By Remark 1 it is enough to show that for each $s_{-i},\in S_{-i}$, for every simple generalized self-odds ratio, equation (15) holds. Without loss of generality, we can assume that the simple generalized self-odds ratio is not infinite (because an infinite one can be inverted and becomes zero), so that no denominator in equation (15) is zero. For each $m\in\left\{2,...,n\right\}$, by consistency, equation (12) for $E=\left\{s_{-i}^{m}\right\}$ and $\left(h_{i}^{1},...,h_{i}^{n}\right)=(h_{i}^{m-1},h_{i}^{m})$ yields
@@
Then, substituting in (15), we obtain equation (14), which coincides with equation (13) for $E=\left\{s_{-i}\right\}$, and thus holds by consistency.
@@ A.2 Completion of the proofs of the theorems
Proof of Theorem 1:
$\mathbf{1\Rightarrow 2)}$ Fix a non-terminal contingency $h$ and let $h^{1},...,h^{n}$ be the next contingencies. Let $\alpha_{1},...,\alpha_{n}$ denote the weights that verify the definition of Bayes-plausibility. We construct probability distributions $(\widetilde{p}(h^{m}|h,s))_{m=1}^{n}$ of reaching each $h^{m}$ given $h$ and each $s\in S(h)$. Fix $m\in\left\{1,...,n\right\}$. If $\alpha_{m}=0$, let $\widetilde{p}(h^{m}|h,s)=0$ for all $s\in\mathcal{S}$. Suppose now that $\alpha_{m}>0$. For each $s\in S(h)$, let
$\widetilde{p}(h^{m}|h,s)=\left\{\begin{tabular}[]{ll}$\frac{\mu(s|h^{m})\alpha_{m}}{\mu(s|h)}$&$\text{if }s\in\mathrm{supp}\mu(\cdot|h)$\\ $\alpha^{m}$&otherwise.\end{tabular}\right$.
@@
Now fix $m\in\left\{1,...,n\right\}$. If $\alpha_{m}=0$, then
$\sum_{s\in S(h)}\widetilde{p}(h^{m}|h,s)=0.$
If $\alpha_{m}>0$, by Bayes-plausibility
@@
$\frac{\widetilde{p}(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}\widetilde{p}(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=\frac{\frac{\mu(s|h^{m})\alpha_{m}}{\mu(s|h)}\mu(s|h)}{\alpha_{m}\sum_{s^{\prime}\in\mathrm{supp}\mu(\cdot|h)}\mu(s^{\prime}|h^{m})}=\mu(s|h^{m}),\qquad\forall s\in S\left(h\right)$
where the last equality follows from (16) and
$\frac{\widetilde{p}(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}\widetilde{p}(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=0=\mu(s|h^{m})\qquad\forall s\notin S\left(h\right)\text{.}$
Thus, $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$ whenever possible.
$\mathbf{2\Rightarrow 3)}$ Fix a system of bets that the agent accepts. We show that it is not a deterministic Dutch book. Fix an initial contingency $h$. Along the paths $\bar{L}$ that start from $h$ and are consistent with some $s\insupp$$\mu(\cdot|h)$, forward-consistency is equivalent to updating by conditioning a joint belief over states and paths with marginal $\mu(\cdot|h)$ over the states. Thus, the Dutch-book argument for conditioning (Teller, 1973; Lewis, 1999) guarantees that the bookmaker either makes zero profit along every $l\in\bar{L}$, or negative profit along some $\bar{l}\in\bar{L}$, under some state $\bar{s}$ that is consistent with $\bar{l}$. [^21] This applies to betting on the state-path pair, and hence, as a subcase, also to our case of betting on the state only. In the second case, the system is not a deterministic Dutch book, as it violates condition of equation (3) with $s=\bar{s}$ and $l=\bar{l}$. The first case is impossible: since acceptance (given the tie-breaking rule) guarantees that the agent expects a positive payoff under every bet along the paths in $\bar{L}$, one could lower the payouts to the agent, preserve acceptance, and give positive profits to the bookmaker along all the paths, contradicting the Dutch-book argument for conditioning.
Proof of Lemma 1. Suppose that there exist two different GORs between some $s$ and $s^{\prime}$. Then, inverting all the DORs in the second GOR and multiplying by the first GOR, [^22] Since the two GORs are different, they cannot be both infinite or both zero, therefore after inverting one of the two their product is not indeterminate. we obtain a new GOR between $s$ and itself, which is not $1$; write it as
$r:=\prod_{i=1}^{n}o(s^{i-1},s^{i}|h^{i}),$
where $n\geq 2$, $s^{0}=s^{n}=s$, and for each $k\in\left\{1,...,n\right\}$,
$o(s^{k-1},s^{k}|h^{k})=\frac{p(h^{k}|s^{k})}{p(h^{k}|s^{k-1})}\frac{\mu(s^{k-1}|h^{k})}{\mu(s^{k}|h^{k})}.$ (17)
Without loss of generality, we can suppose that:
$s^{0},...,s^{n-1}$ are all distinct, because if $s^{j}=s^{k}$, either $r^{\prime}=o(s^{j},s^{j+1}|h^{j})\cdot...\cdot o(s^{k-1},s^{k}|h^{k})$ is a new GOR between $s^{j}$ and itself that is not $1$, or $r^{\prime\prime}=r/r^{\prime}$ is a new GOR between $s$ and itself that is not $1$;
for each k=i,,n1k=i,...,n-1, $h^{k}\not=h^{k+1}$, because if $h^{k}=h^{k+1}=h$, we have $o(s^{k-1},s^{k}|h^{k})\cdot o(s^{k},s^{k+1}|h^{k+1})=o(s^{k-1},s^{k+1}|h)$ and we can substitute the left-hand side with the right-hand side in the GOR (eliminating state $s^{k}$) — likewise, we can suppose $h^{n}\not=h^{1}$;
$r<1$, because if a GOR is larger than $1$, inverting all the DORs we obtain a GOR smaller than $1$ (still between $s$ and itself). $\blacksquare$
Proof of Proposition 1. Consider the GOR given by the definition of $(r,n)$-not-coherent. We fix a $\delta>0$ and use this GOR to construct a $\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$-robust Dutch book. Fix $\varepsilon>0$ and let $\alpha=\sqrt[n]{\frac{1}{r}}$. Define $g(s^{0}|h^{1})$ and $g(s^{1}|h^{1})$ as:
@@
$g(s^{1}|h^{1})$ $=$ $-g(s^{0}|h^{1})\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon=\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon.$
Recursively, for each $m\in\left\{2,...,n\right\}$ define $g(s^{m-1}|h^{m})$ and $g(s^{m}|h^{m})$ as: [^23] Since $h^{m}\not=h^{m-1}$, we are not redefining $g(s^{m-1}|h^{m-1})$.
$g(s^{m-1}|h^{m})$ $=$ $-g(s^{m-1}|h^{m-1})\frac{p(h^{m-1}|s^{m-1})}{p(h^{m}|s^{m-1})}\sqrt[n]{\frac{1}{r}},$ (18)
$g(s^{m}|h^{m})$ $=$ $-g(s^{m-1}|h^{m})\frac{\mu(s^{m-1}|h^{m})}{\mu(s^{m}|h^{m})}+\varepsilon.$ (19)
Complete the system of bets with $g(\widetilde{s}|h)=0$ wherever $g(\widetilde{s}|h)$ is not already defined.
@@
For each $m\in\left\{1,...,n\right\}$, the agent is willing to accept $g(\cdot|h^{m})$ because
$\mu(s^{m-1}|h^{m})g(s^{m-1}|h^{m})+\mu(s^{m}|h^{m})g(s^{m}|h^{m})$
$=$ $\mu(s^{m-1}|h^{m})g(s^{m-1}|h^{m})-g(s^{m-1}|h^{m})\mu(s^{m-1}|h^{m})+\varepsilon\mu(s^{m}|h^{m})=\varepsilon\mu(s^{m}|h^{m})>0\text{,}$
where $\mu(s^{m}|h^{m})>0$ because $r\not$$\infty$. [^24] If $h^{m}=h^{k}=h$ for some $k\not=m$, the same inequality holds for the other involved states, so $g(\cdot|h)$ is accepted. Thus, the agent is willing to accept $g=(g(\cdot|h))_{h\in\mathcal{H}}$.
Let $\hat{p}$ be such that $\max_{h,h^{\prime}\in\mathcal{H},s\in S}\frac{\hat{p}(h|s)}{\hat{p}(h^{\prime}|s)}\frac{p(h^{\prime}|s)}{p\left(h|s\right)}<\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$. For each $m\in\left\{1,...,n-1\right\}$ (recall that $s^{n}=s^{1}$), since $s^{m}\not=s^{m^{\prime}}$ for every $m^{\prime}\not=m$, we have
$\sum_{h\in\mathcal{H}}\hat{p}(h|s^{m})g(s^{m}|h)$ $=$ $\hat{p}(h^{m}|s^{m})g(s^{m}|h^{m})+\hat{p}(h^{m+1}|s^{m})g(s^{m}|h^{m+1})$
$=$ $\hat{p}(h^{m}|s^{m})g(s^{m}|h^{m})-\hat{p}(h^{m+1}|s^{m})g(s^{m}|h^{m})\frac{p(h^{m}|s^{m})}{p(h^{m+1}|s^{m})}\sqrt[n]{\frac{1}{r}}<0,$
[…]
$p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$ $\leq$ $0\iff g(s|h^{1})+\frac{p(h^{n}|s)}{p(h^{1}|s)}g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)\leq 0$
$\iff$ $g(s|h^{1})+\frac{p(h^{n}|s)/\hat{p}(h^{n}|s)}{p(h^{1}|s)/\hat{p}(h^{1}|s)}\frac{\hat{p}(h^{n}|s)}{\hat{p}(h^{1}|s)}g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)\leq 0$
$\implies$ $g(s|h^{1})+\frac{\hat{p}(h^{n}|s)}{\hat{p}(h^{1}|s)}g(s|h^{n})\leq 0$
$\iff$ $\hat{p}(h^{1}|s)g(s|h^{1})+\hat{p}(h^{n}|s)g(s|h^{n})\leq 0$
$\iff$ $\sum_{h\in\mathcal{H}}\hat{p}(h|s)g(s|h)\leq 0\text{.}$
Note that, for $\varepsilon=0$,
$p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)-p(h^{n}|s)g(s^{n-1}|h^{n})\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)+p(h^{n}|s)g(s^{n-1}|h^{n-1})\frac{p(h^{n-1}|s^{n-1})}{p(h^{n}|s^{n-1})}\sqrt[n]{\frac{1}{r}}\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-1}|h^{n-1})\sqrt[n]{\frac{1}{r}}o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)-\beta p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-1})\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\sqrt[n]{\frac{1}{r}}o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-2})\frac{p(h^{n-2}|s^{n-2})}{p(h^{n-1}|s^{n-2})}\left(\sqrt[n]{\frac{1}{r}}\right)^{2}\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)+p(h^{n-2}|s^{n-2})g(s^{n-2}|h^{n-2})\left(\sqrt[n]{\frac{1}{r}}\right)^{2}o(s^{n-2},s^{n-1}|h^{n-1})o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)-p(h^{1}|s^{1})g(s^{0}|h^{1})\frac{\mu(s|h^{1})}{\mu(s^{1}|h^{1})}\left(\sqrt[n]{\frac{1}{r}}\right)^{n-1}o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-2},s^{n-1}|h^{n-1})o(s^{n-1},s|h)\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
$=$ $-p(h^{1}|s)+p(h^{1}|s)r\left(\sqrt[n]{\frac{1}{r}}\right)^{n-1}\left(\sqrt[n]{\frac{1}{r}}-\delta\right)=-p(h^{1}|s)+p(h^{1}|s)r\left(\sqrt[n]{\frac{1}{r}}\right)^{n}\frac{\left(\sqrt[n]{\frac{1}{r}}-\delta\right)}{\sqrt[n]{\frac{1}{r}}}$
$\leq$ $-p(h^{1}|s)+p(h^{1}|s)\frac{\left(\sqrt[n]{\frac{1}{r}}-\delta\right)}{\sqrt[n]{\frac{1}{r}}}<0\text{,}$